# Reconstructing Intelligible Speech from the Pressure Sensor Data in HVACs

### v1

@@ Abstract
Pressure sensors are an integrated component of modern Heating, Ventilation, and Air Conditioning (HVAC) systems. As these pressure sensors operate within the 0–10 Pa range, support high sampling frequencies of 0.5-2 kHz, and are often placed close to human proximity, they can be used to eavesdrop on confidential conversation, since human speech has a similar audible range of 0-10 Pa and a bandwidth of 4 kHz for intelligible quality. This paper presents WaLi, which reconstructs intelligible speech from the low-resolution and noisy pressure sensor data by providing the following technical contributions: (i) WaLi reconstructs intelligible speech from a minimum of 0.5 kHz sampling frequency of pressure sensors, whereas previous work can only detect hot words/phrases. WaLi uses complex-valued conformer and Complex Global Attention Block (CGAB) to capture inter-phoneme and intra-phoneme dependencies that exist in the low-resolution pressure sensor data. (ii) WaLi handles the transient noise injected from HVAC fans and duct vibrations, by reconstructing both the clean magnitude and phase of the missing frequencies of the low-frequency aliased components. Extensive measurement studies on real-world pressure sensors show an LSD of 1.24 and NISQA-MOS of 1.78 for 0.5 kHz to 8 kHz upsampling. We believe that such levels of accuracy pose a significant threat when viewed from a privacy perspective that has not been addressed before for pressure sensors.
Index Terms:
HVAC, pressure sensor, eavesdropping, complex-valued network, magnitude and phase reconstruction.
1 Introduction
@@
The integration of smart sensors into Heating, Ventilation, and Air Conditioning (HVAC) systems has revolutionized building automation and energy management. Today’s HVAC systems utilize a network of smart sensors to monitor environmental and operational parameters in real-time. However, such a rich ecosystem of smart sensors is often considered a “double-edged sword” since they can be used to leak private information. To understand what level of leakage is appropriate, the key question boils down to: What types of sensors can leak private information from today’s HVAC systems and how much information can be inferred from a given sensor data in HVAC systems?
@@
To answer the above question, we need to point out that pressure sensors are an integrated component of HVAC systems. Pressure sensors often operate within the pressure range of 0-10 Pa and have a high sampling frequency of 0.5-2 kHz [1, 2, 3], which is essential for the dynamic control of fans, dampers, and air handling units for real-time monitoring and fast response in today’s HVAC systems [4, 5, 6]. These pressure sensors are often placed close to humans and installed in room walls, near diffusers, or within ventilation grilles where human occupancy is high to control indoor air quality (IAQ), to monitor thermal comfort, or room-level pressure regulation in spaces like offices, hospitals, or cleanrooms. As these pressure sensors operate within the 0–10 Pa range, support high sampling frequencies of 0.5-2 kHz, and are often placed close to human proximity, they can be used to eavesdrop on confidential conversation, since human speech has a similar audible range of 0-10 Pa [7] and a bandwidth of 4 kHz for intelligible quality.
@@
To eavesdrop on a confidential conversation with intelligible quality, the attacker must reconstruct speech from pressure sensor data with unrestricted vocabulary by answering the following two questions: (1) Will it be possible to capture intelligible speech of 4 kHz bandwidth with 0.5-2 kHz sampling frequencies using pressure sensors? And (2) How do we handle transient noise injected from HVAC fans, duct vibrations, physical shocks, or turbulent airflow while reconstructing intelligible speech from pressure sensor data located in HVAC systems?
@@
In this paper, we answer the above two questions by proposing WaLi (Wall can Listen), which can recover intelligible speech by reconstructing bandwidth up to 4 kHz from a lowest sampling frequency of 500 Hz from pressure sensor in transient noisy conditions. WaLi employs the following two technical strategies to answer the above two questions:
@@
Strategy 1. With a sampling rate of 0.5-2 kHz, high-frequency speech components are severely aliased and truncated in pressure sensors, while a few lower pitch frequencies are preserved. Unfortunately, a few low pitches are not sufficient to provide perfect intelligibility [8] as most of the formants at high frequencies will be missing from the bandwidth. To reconstruct intelligible speech from the severely aliased spectrogram of pressure sensors, the missing high frequencies should be reconstructed from the low-frequency aliased components. To make this happen, WaLi employs Conformers [9], combining the strengths of Convolutional Neural Networks (CNNs) and Transformers to capture local and global dependencies between low-frequency pitches and missing high frequencies. In addition to conformers, WaLi employs Complex Global Attention Block (CGAB) to capture the long-range inter-phoneme correlations that exist among pitches and harmonics along both the time and the frequency axes in a spectrogram. Prior work [10] published in USENIX only captures correlations along the time axis.
@@
Strategy 2. To prevent the transient noise of HVAC from impacting speech reconstruction, WaLi reconstructs both the clean magnitude and phase of the missing frequencies from the low-frequency aliased components. To jointly reconstruct clean magnitude and phase, WaLi is designed as a complex-valued network, which can enhance both the amplitude and phase of pressure sensor data using complex-valued time-frequency (T-F) spectrogram [11]. As complex-valued T-F spectrograms have both magnitude and phase information, WaLi takes the complex-valued T-F spectrograms as input and removes noisy phases from them, motivated by the fact that phase plays a crucial role in speech enhancement [12]. Prior works [10, 13, 14] only use real-valued T-F spectrograms. Therefore, they cannot reconstruct clean phase under noisy conditions. WaLi employs complex multi-resolution STFT loss to reconstruct clean magnitude and phase from the noisy pressure sensor data.
@@
We extensively evaluate the effectiveness of WaLi through real-world experiments, using five evaluation metrics - LSD, NISQA-MOS, PESQ, STOI, and SI-SDR. The full form of these terms is given in Section 7.2. Our findings highlight the serious privacy implications of pressure sensors in HVAC systems. The implications could be particularly severe in critical environments, such as industrial facilities, corporate offices, and healthcare institutions, where eavesdropping can expose sensitive information and compromise privacy. Our main contributions are summarized as follows:
@@
2. We use complex-valued architecture to jointly reconstruct both the clean magnitude and phase from the noisy pressure sensor data and extensively evaluate the effectiveness of WaLi using five evaluation metrics - LSD, NISQA-MOS, PESQ, STOI, and SI-SDR. A demonstration of the reconstructed audio is provided in the following link: WaLi.
2 Related Work
Acoustic eavesdropping: Acoustic eavesdropping using different sensors in different modalities is extensively explored in the literature. For example, lasers [15, 16], inertial measurement units (IMU) [17, 18, 19, 20, 13], wireless signals [21, 22, 23, 24, 25, 26], optical sensors [27, 28], vibration motors [29], and hard drives [30] are explored to reveal great threats to speech privacy. mmEcho [21], mmEve [31], and mmSpy [32] use mmWave sensors to capture vibrations and reconstruct speech from vibration data. However, these works cannot reconstruct speech with full intelligibility from the narrowband vibration data. IMU sensors [18, 19] and gyroscopes [20] are used to eavesdrop using the vibration induced by sound in IMUs and gyroscopes. They can do digit inference, gender recognition, and limited hot-word reconstruction. However, these works still suffer from the narrowband condition of vibration-based side channels and can only recover band-limited sound with damaged speech intelligibility. GlowWorm [33] and Lamphone [28] employ optical sensors to record vibrations and are used to recover speech with limited intelligibility. However, they do not recover clean phases, and, as a result, they will not work in transient noise conditions.
Recently, AccEar [13] and VibSpeech [10] revealed the possibility of using generative adversarial networks to reconstruct intelligible speech with unrestricted vocabulary by learning the mapping between low-frequency pitches and missing high frequencies. Our proposed WaLi solves the following two problems of AccEar [13] and VibSpeech [10]:
1. AccEar [13] assumes that abundant ground-truth speech from the target victim is available for training, which is not realistic. Compared to AccEar [13], our WaLi does not require any {audio, pressure sensor data} pair from a specific target victim for training to recover intelligible audio. This makes our attack model realistic, as it is not practical to have access to the victim’s speech for training.
2. Compared to VibSpeech [10], our WaLi can reconstruct intelligible audio in the presence of transient noise injected from HVAC fans, duct vibrations, physical shocks, or turbulent airflow. VibSpeech [10] has not been tested in the presence of a such transient noise. Moreover, our WaLi does not require SpkEnc-type encoders and extra vocoders like VibSpeech, as our WaLi use complex-valued network to extract features from magnitude and phase of the speech.
Pressure sensor-based eavesdropping: Recently, BaroVox [34] proposed to use pressure sensors to eavesdrop on speech in the cleanroom. The main differences between WaLi and BaroVox are: (1) BaroVox only recognizes hot words or phrases, whereas WaLi can reconstruct intelligible speech with unrestricted vocabulary. (2) BaroVox has not been tested for the transient noise condition. Therefore, it is not realistic for the real-world scenario in HVAC systems.
3 Background
3.1 Human Voice and Its Pressure Range
The human voice generates sound by producing acoustic waves that propagate through the air. These waves create fluctuations in sound pressure, typically measured in Pascals (Pa). A comparative study of the voice pressure range is given in Table I for different human voice conditions.
Condition in dB in Pa
Quiet whisper 40 dB 0.02 Pa
Conversation 60 dB 0.2 Pa
Loud speech 70 dB 0.6 Pa
Shouting 100-115 dB 6-10 Pa
TABLE I: Pressure ranges for different voice conditions.
@@ 3.1 Human Voice and Its Pressure Range
Table I indicates that the pressure of speech signals is typically restricted between 0 to 10 Pa for a wide range of intelligible speech. Therefore, to successfully eavesdrop on speech signals, pressure sensors should also need to be sensitive in the same range of 0 to 10 Pa.
3.2 Pitch, Phonemes, and Intelligible Bandwidth
3.3 Pressure Sensor in HVACs
Differential pressure sensors (DPSs) are the state-of-the-art (SOTA) sensors for HVAC systems due to their better control, accurate measurement, and reliable operations [35]. Hence, in the rest of the paper, we use the DPS and the pressure sensor terms interchangeably. DPSs measure the difference between any two pressure levels.
DPSs have an elastic diaphragm, which collects the input pressure from the environment. The elastic diaphragm is placed between two pressure input ports $P_{1}$ and $P_{2}$ (see Fig. 1(Left)). The diaphragm senses the differential pressure $P_{1}$ $P_{2}$ applied to the pressure input ports by changing its shape. The change in the shape of the diaphragm is converted to a proportional output voltage by using a transducer and a Wheatstone bridge. The diaphragm is sensitive to acoustic pressure and can pick up sound pressure when someone speaks. In real-world HVAC systems, sampling tubes are connected with either one or both input ports of DPSs. The other end of the sampling tube is connected to pressure pickup devices (see Fig. 1 (Right)), which senses pressure in the environment.
@@ 3.3 Pressure Sensor in HVACs
Figure 1: (Left) Internals of a DPS. (Right) The accessories connected with DPSs in a real-world HVAC system.
@@
3.4 Pressure Range and Sampling Frequencies of Differential Pressure Sensors (DPSs) in HVACs
Sections 3.1 and 3.2 point out that for eavesdropping using DPSs in HVACs, the pressure sensors should be sensitive within 0-10 Pa and have a sampling frequency close to 8 kHz. In HVAC systems, DPSs operating within the 0–10 Pa range and supporting high sampling frequencies within 0.5-2 kHz are typically used to achieve precise environmental control and energy efficiency. These sensors are used in a variety of applications where small pressure changes must be detected accurately and quickly. Common use cases include air filter monitoring, cleanroom pressure regulation, static duct pressure control, and variable air volume (VAV) system management. High-frequency sampling enables real-time responsiveness to transient pressure fluctuations, which is essential for dynamic control of fans, dampers, and air handling units. By providing fast and accurate data, these sensors support improved indoor air quality, occupant comfort, and operational efficiency. A summary of the pressure sensor range and sampling frequencies in HVACs is given in Table II. Table II shows that pressure sensors in HVACs are sensitive to the audible pressure range of 0-10 Pa.
@@ 3.4 Pressure Range and Sampling Frequencies of Differential Pressure Sensors (DPSs) in HVACs
TABLE II: Human audible pressure range and sampling frequencies of pressure sensors in HVACs.
3.5 Impact of Speech on Pressure Sensors
Table II indicates that DPSs have a sampling frequency of 0.5-2 kHz. However, at least 8 kHz sampling is required to recover intelligible speech. Therefore, a natural question is: Will it be possible to capture intelligible speech with at least 0.5 kHz sampling frequency using pressure sensors?. The pitch frequencies for a male speaker vary between 85-180 Hz, whereas for a female speaker, they vary between 165-255 Hz [39]. Although the 0.5 kHz sampling rate can capture pitches and a few harmonics, it is not sufficient to provide perfect intelligibility [8], as most harmonics at high frequencies will be missing from the bandwidth.
@@ 3.5 Impact of Speech on Pressure Sensors
Figure 3: (Left) Clean speech in the absence of noise. (Right) Noisy speech in the presence of transient noise.
3.6 Transient Noise and Phase Reconstruction
DPSs operating within a low-pressure range of 0–200 Pa and at high sampling frequencies of 0.5-2 kHz, the impact of noise on DPSs becomes particularly significant. The presence of transient noise sources, such as HVAC fans, duct vibrations, physical shocks, or turbulent airflow, injects substantial noise into DPSs in the form of pressure spikes and fluctuating readings. The noisy readings from DPSs impact the intelligible speech reconstruction for successful eavesdropping. Although earlier speech reconstruction methods [13, 10, 34] focus on estimating the magnitude spectrum of missing higher frequencies, the phase spectrum is equally important [12] to produce natural and intelligible audio, particularly under transient noise conditions in real-world HVAC systems. Fig. 3 shows the impact of noise on pressure sensor data at 0.5 kHz sampling frequency. Transient noise corrupts the phase and magnitude of the speech signal in HVAC systems and change the Signal-to-Noise Ratio (SNR) from 8 dB to 3.5 dB.
Application Area Sensor Location Purpose
Differential control [40] Wall and ceiling near doorways Maintain pressurization in isolation rooms in hospitals
Controlled ventilation [41] Inside duct or vent near occupants Adjust airflow based on occupancy level
Lab pressure monitor [42] Inside room or partition walls Protect from contamination by maintaining air barriers
Residential IAQ [43] Return air ducts in living areas Balance pressure and airflow for occupant comfort
TABLE III: DPSs placed close to humans can do eavesdropping.
4 Pilot Study
Here, we discuss the feasibility of eavesdropping on intelligible speech using DPSs in HVAC systems.
Proximity to sound sources and humans: DPSs are an integrated component of today’s HVAC systems. However, to eavesdrop using DPSs, the DPSs should be close to humans or a sound source; otherwise, the feasibility of the acoustic eavesdropping using DPSs will be reduced. We identify a few locations in HVACs, summarized in Table III, where DPSs are placed close to humans.
Table III indicates that DPSs placed close to humans are typically used for indoor air quality (IAQ) control, thermal comfort monitoring, or room-level pressure regulation in spaces like offices, hospitals, or cleanrooms. These sensors operate in low-pressure ranges (e.g., 0–250 Pa) and are often installed in room walls, near diffusers, or within ventilation grilles where human occupancy is high. To support this claim, we survey an anonymous cleanroom in our educational premises and point out the location of DPSs near the human occupancy that is shown in Fig. 4.
Figure 4: Pressure ports are located at the hallway entrance and inside rooms of a lab facility in human proximity.
Fig. 4 supports the fact that DPSs can be found in room entrances and inside rooms in practical HVAC systems in human proximity. Therefore, they can be a source of eavesdropping. The main challenges are that DPSs are noisy and use a sampling frequency of 0.5-2 kHz (refer to Table II). Therefore, we propose WaLi to recover full intelligible speech from the aliased pressure sensor data by reconstructing both amplitude and phase of the speech signal.
A pilot experiment: We prepare a testbed as a pilot study to support our idea of eavesdropping using pressure sensors with an unrestricted vocabulary. We use an industry-used DPS from Setra with part# 26410R1WD11T1E [44]. It has two input ports. We connect two vinyl sampling tubes with inner diameters of 3/16” and 5/16” [45], a pressure pickup device with part# A-417A [46] with the two input ports (see Fig. 1). A volunteer speaks from 0.5 m distance from the pressure pickup device. We record the output data from the DPS sensor using an oscilloscope with a sampling frequency of 1 kHz. The recorded data from the DPS are shown in Fig. 5 with its matching ground-truth audio. The ground truth audio is sampled at 8 kHz.
@@ 4 Pilot Study
Figure 5: Victim is talking close to the pressure pickup device in his room.
In Fig. 5, the time domain waveform indicates a strong correlation between the ground truth audio and the pressure sensor data. We measure the intelligibility of the pressure sensor data by taking the ground truth as a reference using a metric named Perceptual Evaluation of Speech Quality (PESQ). PESQ is widely used to evaluate how a degraded speech signal compares to a reference (ground truth) signal, closely modeling human perception. The PESQ for the pressure data is 0.98. PESQ can have values between -0.5 to 4.5, where -0.5 means low perceptual quality and 4.5 means perceptually close to the ground truth. The 0.98 of PESQ for the pressure data indicates that the DPS captures low-grade perceptual audio, which is not intelligible (to be intelligible PESQ value of more than 1.4 is required). Therefore, it indicates that DPS can be used to eavesdrop by an attacker if the attacker can reconstruct intelligible audio from the severely aliased and low-grade perceptual data of DPSs.
5 Threat Model
We discuss the components of the threat model below.

### v2

# WaLi: Can Pressure Sensors in HVAC Systems Capture Human Speech?

@@ Abstract
Pressure sensors are an integrated component of modern Heating, Ventilation, and Air Conditioning (HVAC) systems. As these pressure sensors operate within the 0–10 Pa range, support high sampling frequencies of 0.5-2 kHz, and are often placed close to human proximity, they can be used to eavesdrop on confidential speech, since human speech has a similar audible range of 0-10 Pa and a bandwidth of 4 kHz for intelligible quality. This paper presents WaLi, which reconstructs intelligible speech from the low-resolution and noisy pressure sensor data with the following technical contributions: (i) WaLi reconstructs intelligible speech from a minimum of 0.5 kHz sampling frequency of pressure sensors, whereas previous work can only detect hot words/phrases. WaLi uses a complex-valued conformer and Complex Global Attention Block (CGAB) to capture inter-phoneme and intra-phoneme dependencies that exist in the low-resolution pressure sensor data. (ii) WaLi handles the transient noise injected from HVAC fans and duct vibrations by reconstructing both the clean magnitude and phase of the missing frequencies of the low-frequency aliased components. We evaluate our attack on practical HVAC systems located in two anonymous industrial facilities. Extensive studies on real-world pressure sensors show an LSD of 1.24 and a NISQA-MOS of 1.78 for 0.5 kHz to 8 kHz upsampling. We believe that such levels of accuracy pose a significant threat when viewed from a privacy perspective that has not been addressed before for pressure sensors. We also provide defenses for the attack.
I Introduction
@@
The integration of smart sensors into Heating, Ventilation, and Air Conditioning (HVAC) systems has revolutionized building automation and energy management. However, such a rich ecosystem of smart sensors is often considered a “double-edged sword” since they can be used to hamper privacy. To understand what level of leakage is appropriate, the key question boils down to: What types of sensors can leak privacy from today’s HVAC systems and how much information can be inferred from a given sensor data in HVACs?
@@
To answer the above question, we need to point out that pressure sensors are an integrated component of HVAC systems. Pressure sensors often operate within the pressure range of 0-10 Pa and have a high sampling frequency of 0.5-2 kHz [1, 2, 3], which is essential for the dynamic control of fans, dampers, and air handling units for real-time monitoring and fast response in today’s HVAC systems [4, 5, 6]. These pressure sensors are often placed close to humans and installed in room walls, near diffusers, or within ventilation grilles where human occupancy is high to control indoor air quality (IAQ), to monitor thermal comfort, or room-level pressure regulation in spaces like industries, hospitals, or cleanrooms. As these pressure sensors operate within the 0–10 Pa range, support high sampling frequencies of 0.5-2 kHz, and are often placed close to human proximity, they can eavesdrop on confidential speech, since human speech has a similar audible range of 0-10 Pa [7] and a bandwidth of 4 kHz for intelligible quality.
@@
To eavesdrop on confidential speech with intelligible quality, the attacker must reconstruct speech from pressure sensor data by answering the following two questions: (1) How to reconstruct intelligible speech of 4 kHz bandwidth from 0.5-2 kHz sampling frequencies of pressure sensors? And (2) How do we handle transient noise injected from HVAC fans, duct vibrations, physical shocks, or turbulent airflow while reconstructing intelligible speech from pressure sensor data?
@@
In this paper, we answer the above two questions by proposing WaLi (Wall can Listen), which can reconstruct intelligible speech with a bandwidth up to 4 kHz from a lowest sampling frequency of 500 Hz of pressure sensor in transient noisy conditions. WaLi employs the following two technical strategies to answer the above two questions:
@@
Strategy 1. With a sampling rate of 0.5-2 kHz, high-frequency speech components are severely aliased and truncated in pressure sensors, while a few lower pitch frequencies are preserved. Unfortunately, a few low pitches are not sufficient to provide perfect intelligibility [8] as most of the formants at high frequencies will be missing from the bandwidth. To reconstruct intelligible speech from the severely aliased spectrogram of pressure sensors, the missing high frequencies should be reconstructed from the low-frequency aliased components. To make this happen, WaLi employs conformers [9], combining the strengths of Convolutional Neural Networks (CNNs) and Transformers to capture local and global dependencies between low-frequency pitches and missing high frequencies. In addition to conformers, we design Complex Global Attention Block (CGAB) to capture the long-range inter-phoneme correlations that exist among pitches and harmonics along both the time and the frequency axes in a spectrogram. Prior work [10] published in USENIX only captures correlations along the time axis.
@@
Strategy 2. To prevent the transient noise of HVAC from impacting speech reconstruction, WaLi reconstructs both the clean magnitude and phase of the missing frequencies from the low-frequency aliased components. To jointly reconstruct clean magnitude and phase, WaLi is designed as a complex-valued network, which can enhance both the amplitude and phase of pressure sensor data using complex-valued time-frequency (T-F) spectrogram [11]. As complex-valued T-F spectrograms have both magnitude and phase information, WaLi takes the complex-valued T-F spectrograms as input and removes noisy phases from them, motivated by the fact that phase plays a crucial role in speech enhancement [12]. Prior works [10, 13, 14] only use real-valued T-F spectrograms and cannot reconstruct a clean phase under noisy conditions. Moreover, WaLi employs a complex multi-resolution Short-Time Fourier Transformation (STFT) loss to reconstruct clean magnitude and phase from the noisy pressure sensor data.
@@
We extensively evaluate the effectiveness of WaLi in two real-world anonymous industrial facilities, using six evaluation metrics - LSD, NISQA-MOS, PESQ, STOI, WER, and SI-SDR (see Section VI-B for the full form of these terms). Our findings highlight the serious privacy implications of pressure sensors in HVAC systems. The implications could be particularly severe in critical environments, such as industrial facilities, corporate offices, and healthcare institutions, where eavesdropping can expose sensitive private information. Our main contributions are summarized as follows:
@@
2. We use complex-valued architecture to jointly reconstruct both the clean magnitude and phase from the noisy pressure sensor data and extensively evaluate the effectiveness of WaLi using six evaluation metrics - LSD, NISQA-MOS, PESQ, STOI, WER, and SI-SDR. A real-world demonstration of the reconstructed audio is provided in the following link: WaLi.
II Background
II-A Human Voice’s Pressure Range and Intelligible Bandwidth
@@
The human voice creates fluctuations in sound pressure, measured in Pascals (Pa). A comparative study of the voice pressure range, given in Table XI of Appendix A, indicates that pressure of speech signals is typically restricted between 0-10 Pa for a wide range of conditions [15]. Therefore, to successfully eavesdrop on speech, pressure sensors should also need to be sensitive in the same range of 0-10 Pa. Moreover, for speech intelligibility, only 8 kHz sampling is sufficient, as up to 4 kHz bandwidth is enough for intelligible speech.
@@
Figure 1: (Left) Internals of a DPS. (Right) The accessories connected with DPSs (part# SETRA264) [16] in a real-world HVAC system.
II-B Pressure Sensor in HVACs
Differential pressure sensors (DPSs) are the state-of-the-art sensors for HVAC systems due to their accurate measurement and reliable operations [17]. Hence, in the rest of the paper, we use the DPS and the pressure sensor terms interchangeably.
DPSs have an elastic diaphragm placed between two pressure input ports $P_{1}$ and $P_{2}$ (see Fig. 1 (Left)). The diaphragm senses the differential pressure $P_{1}$ - $P_{2}$ applied to the input ports by changing its shape. The change in the shape of the diaphragm is converted to a proportional output voltage by using a transducer and a Wheatstone bridge [18]. The diaphragm is sensitive to acoustic pressure and can pick up sound pressure when someone speaks. In real-world HVAC systems, like part# SETRA264 [16], sampling tubes are connected with either one or both input ports of DPSs. The other end of the sampling tube is connected to pressure pickup devices (see Fig. 1 (Right)), which senses pressure in the environment.
@@ II-B Pressure Sensor in HVACs
TABLE I: Human audible pressure range and sampling frequencies of pressure sensors in HVACs.
@@
II-C Pressure Range and Sampling Frequencies of DPSs
Sections II-A points out that for eavesdropping using DPSs in HVACs, the DPSs should be sensitive within 0-10 Pa and should have a sampling frequency close to 8 kHz. In HVAC systems, DPSs operating within the 0–10 Pa range and supporting high sampling frequencies within 0.5-2 kHz are typically used to achieve precise environmental control where small pressure changes must be detected accurately and quickly. High-frequency sampling enables real-time responsiveness to transient pressure fluctuations, which is essential for dynamic control of fans, dampers, and air handling units. Common use cases include air filter monitoring [20], cleanroom pressure regulation [22], static duct pressure control, and variable air volume (VAV) management [21]. A summary of the pressure sensor range and sampling frequencies in HVACs is given in Table I, which shows that DPSs in HVACs are sensitive to the audible pressure range of 0-10 Pa.
@@ II-C Pressure Range and Sampling Frequencies of DPSs
Figure 2: (Left) Clean speech captured by SDP 1108 in the absence of noise. (Right) Noisy speech captured by SDP 1108 in the presence of transient noise.
@@
Figure 3: (Left) A brief overview of the threat model - WaLi. The victim is unknowingly talking close to the pressure pickup device.
II-D Transient Noise and Importance of Phase Reconstruction
The presence of transient noise sources, such as HVAC fans, duct vibrations, physical shocks, or turbulent airflow, injects substantial noise into DPSs in the form of pressure spikes and fluctuating readings [23]. The impact of noise becomes particularly significant within a low-pressure range of 0–10 Pa and at high sampling frequencies of 0.5-2 kHz. The noisy readings from DPSs impact the intelligible speech reconstruction for successful eavesdropping. Although earlier speech reconstruction methods [13, 10, 24] focus on estimating the magnitude spectrum of missing higher frequencies, the phase spectrum is equally important [12] to produce natural and intelligible audio, particularly under transient noisy conditions in real-world HVAC systems. Fig. 2 shows the impact of noise on pressure sensor data at 0.5 kHz sampling frequency. Transient noise corrupts the phase and magnitude of the speech signal collected by a DPS (SDP 1108 [21]) and deteriorates the Signal-to-Noise Ratio (SNR) from 8 dB to 3.5 dB.
III Threat Model
We discuss the threat model of WaLi below (see Fig. 3).
a) Proximity to sound sources and humans: To prove that DPSs are often located close to humans, we have evaluated two anonymous facilities - one is an industrial facility and the other is an FDA-compliant cleanroom (see Fig. 4).
Figure 4: (Left & Middle) Pressure ports are located at hallway entrance and corridor of a cleanroom, and (Right) inside rooms of an industrial facility.
Fig. 4 supports the fact that DPSs are often found in room entrances, corridors, inside rooms near diffusers, etc., in practical HVAC systems in human proximity and they operate in low-pressure ranges (e.g., 0–250 Pa). Therefore, they can be a source of eavesdropping. The main challenges are that DPSs are noisy and use a sampling frequency of 0.5-2 kHz (refer to Table I). Therefore, we propose WaLi to recover full intelligible speech from the aliased pressure sensor data by reconstructing both amplitude and phase of the speech signal.